# EnvGuard : Guaranteeing Environment-Centric Properties in Web of Things

### 06 Dec 2023 v1

@@ Abstract.
With the accelerated advancement of IoT, diverse devices are ubiquitously deployed in environments. Building on this, Web of Things (WoT) further integrates fragmented device services and provides unified interfaces using standardized Web technologies, promoting the development and deployment of WoT applications to sense and regulate the environment. However, disparate WoT applications independently control devices in the WoT environment, causing interference among devices and with the environment. This results in device behaviors that deviate from user expectations, causing violations of the user’s desired environment properties. The intricate interplay of applications, user activities, and environment changes makes identifying and resolving potential violations a complex task. In this paper, we introduce EnvGuard, an environment-centric approach for property description, violation identification, and resolution in WoT environment. EnvGuard proposes a conceptual schema to model the relationship between device services and environment context, and automatically extends the conceptual schema into a specific environment representation based on device and space information. Furthermore, EnvGuard employs a template-based approach, enabling users to describe spatial and temporal properties based on the abstract device effects on the environment, and translating properties description into formal expressions. EnvGuard adopts a hybrid model checking method to respectively identify the spatial and temporal violations, and a resolution strategy that align with user intention is proposed to resolve violations. We evaluate EnvGuard through user studies and our proposed dataset, which is constructed by collecting real-world data from a laboratory WoT environment and manually labeling ten types of violations. The results confirm the usability, feasibility and efficiency of EnvGuard.
WoT environment, modeling, violation identification, dataset
@@ 1. Introduction
The vigorous development of Internet of Things (IoT) has accelerated the deployment of smart devices in human living environments. This prosperity has yet caused fragmentation in IoT landscape (Kojima and Sakazawa, 2021), making interoperability between different devices difficult (Lu et al., 2023; Sharma et al., 2017). To this end, Web of Things (WoT) is proposed as a paradigm complementary to IoT, which standardizes the device sensing and actuating capabilities as perceptual event and executable action services through leveraging Web technologies (Sun et al., 2023; Zyrianoff et al., 2022).
Building on the paradigm, various WoT applications can be developed and deployed in WoT environments (Guinard et al., 2011) for automation requirements. In a typical smart office, a humidity control application can automatically open the window when indoor humidity is high. Another more advanced smart meeting application can automate meeting room preparation by coordinating devices like printers, lights, and air conditioners.
Despite the convenience, several studies highlight that the runtime behaviors of devices may often diverge from user expectations (Pradeep et al., 2021; Huang et al., 2021; Chaki et al., 2020; Chaki and Bouguettaya, 2020; Wang et al., 2019; Shah et al., 2019), causing violations of the user’s desired environment properties, i.e., constraints which should always be satisfied by the WoT environment. Multiple factors contribute to the unintended behaviors.
Applications may have inherent flaws if developers overlook the interaction between the complex environment context and context-sensitive device services (Manca et al., 2019; Corno et al., 2019; Zhao et al., 2020), leading to unintended behaviors. For example, the humidity control application triggers a window-opening action due to high indoor humidity during rain, which is counterproductive due to the high outdoor humidity and poses a risk damaging indoor items with rainwater. Additionally, violations may arise between WoT applications due to the lack of coordination between them (McCall et al., 2021; Zhang et al., 2022; McCall et al., 2023). For example, the meeting application activates an air conditioner to cool the environment before a meeting, while the humidity control application opens the window to ventilate the room, resulting in a waste of energy and a poor user experience. Unpredictable environment changes and user activities can also induce violations (Chaki et al., 2020; Camacho, 2014). For example, a user might manually open the meeting room window on a sunny day and leave, followed by the activation of an air conditioner by the meeting application, and subsequently, a change in weather to rain.
In the WoT environment, the interplay of applications (Zhao et al., 2021), user activities (Chaki and Bouguettaya, 2021a), and environmental changes (Zhang et al., 2019) complicates the identification and resolution of potential violations, even for experts. To this end, existing approaches support guaranteeing properties preset internally  (Palekar et al., 2019; Trimananda et al., 2020; Bu et al., 2023) or specified explicitly by end users (Zhang et al., 2019; Bu et al., 2018), conducting automated verification of property violations caused by WoT applications (Yagita et al., 2015; Mohsin et al., 2016), and adopting an intercepting or revoking strategy to repair the application (Chaki and Bouguettaya, 2021b; Nandi and Ernst, 2016). However, challenges still remain in many aspects.
@@
The interplay between environment context and the context-sensitive device services implies that existing violation detection methods for applications may overlook their intricate interrelation. For example, opening and closing a window can have varying effects on indoor temperature, humidity, and air quality under different outdoor conditions. Identifying violations only based on the window’s state can lead to omissions or misjudgments.
@@
The definition of property fails to meet the customization requirements when it is predefined and built-in, while accurately specifying properties by configuring related device services and environment context is challenging. For example, to avoid conflicts between heating and cooling devices, users must consider all devices potentially affecting temperature (e.g., air conditioner, window, and heater) and the related environment context (e.g., weather, outdoor temperature).
@@
Environment properties span spatial and temporal aspects, while existing methods primarily address spatial environment context, neglecting the temporal trace of device services. For instance, ensuring energy conservation by detecting prolonged lighting in empty rooms requires a temporal analysis of human detection events and light-switch actions.
@@
The intercepting or revoking resolve strategy may be inconsistent with user intention, as the violation actions possibly contain user requirements. For example, if a user opens a window while raining, closing it as a resolution might be contrary to the ventilation intent.
To address these challenges, we propose EnvGuard as shown in Figure 1, an environment-centric approach that offers expressive descriptions and comprehensive guarantees for properties through an environment-centric perspective. EnvGuard first provides a conceptual schema to model the interplay between environment context and context-sensitive device services, and automatically generates a representation of a WoT environment based on device and space information. Subsequently, EnvGuard defines eight templates to enable end-users to specify spatial and temporal environment properties through the effects of device services, which can be translated into LTL (Linear Temporal Logic) and MTL (Metric Temporal Logic). During offline phase, EnvGuard separately analyzes the LTL describing spatial state properties and the MTL describing temporal trajectory properties by adopting a hybrid model-checking technique, identifying potential violations and their possible resolution actions. Through a user study on user preferences for resolving different types of violations, a resolution strategy is proposed to ensure the consistency of resolution action and user’s intentions. At runtime, EnvGuard checks spatial state and temporal trace property in an event-driven manner and, based on the resolution strategy, selects non-conflicting actions to resolve identified violations without introducing new ones.
@@
To evaluate EnvGuard, we construct a dataset by collecting real-world data from a laboratory WoT environment and manually labeling ten types of violations. We implement EnvGuard as a prototype and initialize it in the laboratory environment. The usability of the provided GUI tool is evaluated by conducting a second user study, and the performance for the violation identification, resolve, and runtime efficiency is evaluated by the dataset, which is designed to support in-depth analysis of the violations in the WoT environment.
We summarize our contributions as follows:
An environment modeling and representation method to describe the interplay between environment context and device services.
An expressive property description method to assist users in effortlessly defining spatial and temporal properties.
An environment-centric approach for identifying and resolving the violations of user-defined properties in the WoT environment in consideration of the interplay between environment context and the context-sensitive device services.
A dataset for supporting environment-level violation identification, which is publicly available(EnvGuard, 2023) for encouraging further research.
@@ 2. Environment Modeling
In this section, we present a schema modeling the interplay between environment context and device services, and detail its automatic construction from device and space data.
@@ 2.1. Conceptual Schema of Environment
The conceptual schema of the environment is illustrated in Figure 2, describing the concepts and the relationship of the spaces within the environment, the events and actions provided by devices, and the effects that actions can produce on the environment under the specific preconditions. We further formalize each concept in the conceptual schema.
@@
Space describes the spaces within the environment, which is formally defined as a tuple of ¡id, type, Attributes, States¿. id is a unique identifier of the space and type is the category of the space (e.g., meeting room, corridor, etc.). Attributes are a set of space attributes defined as a map $\{(a_{1}:v_{1}),...,(a_{n}:v_{n})\}$ where each $a_{i}$ is the attribute name (e.g., area) and $v_{i}$ is the corresponding value (e.g., 50 $m^{2}$). And States is a set of variables of the space defined as a map $\{(s_{1}:v_{1}),...,(s_{n}:v_{n})\}$ where each $s_{i}$ is the state name (e.g., temperature) and $v_{i}$ is the corresponding value. The continuous state values are discretized (e.g., discretize the temperature into high, medium, and low). The topological structure of different spaces is described by the ”adjacentTo” and ”reachableTo” relationship.
@@
Device is represented as a tuple of ¡id, type, state¿, indicating the unique identifier, category (e.g., air conditioner, window), and current state (e.g., on, off). The ”locatedIn” relationship specifies the space where a device is deployed.
@@
Event is one type of the device service describing the environmental state change which is provided by sensor. Event is defined as a tuple of ¡event_type, location, target_state, payload, Topic¿. event_type indicates the category of the event (e.g., Temperature_High), location indicates the space where the event occurs, target_state specifies the space state impacted by the event (e.g., temperature), and payload is the current value of the changed state (e.g., high). Topic is defined as a map $\{(id_{1}:t_{1}),...,(id_{n}:t_{n})\}$ where each $id_{i}$ indicates the id of device that can provide the event, and each $t_{i}$ indicates the message queue topic for subscribing to the event provided by the device. The ”providedBy” relationship specifies the device providing the event, and the ”update” relationship specifies the target space where the state change has occurred.
@@
Action is the other type of device service provided by actuator, which is defined as a tuple of ¡action_type, location, target_state, payload, URL¿. action_type denotes the category of the action (e.g., AC_TurnOn), location indicates where the action is executed, target_state specifies the device whose state is impacted by the action, and the payload is the value of the current device state. URL is defined as a map $\{(id_{1}:url_{1}),...,(id_{n}:url_{n})\}$ where each $id_{i}$ indicates the id of the device that can provide the action, and each $url_{i}$ indicates the API address for executing the action provided by the device service. The ”providedBy” relationship specifies the device performing the action, and the ”has” relationship specifies the effect that the action execution will cause on the environmental contexts.
@@
Effect describes the impact of an action on the environment state, which is defined as a pair of ¡state_name, effect_type¿. state_name specifies the space state impacted by the action (e.g., temperature), and effect_type specifies the effect of the action on the target_state (e.g., increase, decrease, etc.). The ”affect” relationship indicates the space impacted by this effect, and the ”constrainedBy” relationship specifies the prerequisite conditions, PreCondition, that need to be satisfied for the effect to take place. PreCondition is defined as a tuple of ¡state_name, operator, target_value¿, where state_name specifies the space state that needs to be satisfied, operator specifies the comparison operator (e.g., ¿, ¡, etc.), and target_value specifies the value that needs to be satisfied. Specifically, the target_value can be set either by a specific value or referring to a space state. For example, the target_value can be set to high or refer to the temperature of the adjacent space. The ”referTo” relationship specifies the space that the state_name and target_value refer to.
@@ 2.2. Environment Representation Construction
All instances generated by the device description format can be associated with space instances based on the device’s location information (e.g., the window instance is connected to the Lab01 instance through the ”locatedIn” relationship), completing the construction of environment representation. The complete algorithm for constructing environment representation is presented in A.1.
@@ 3. Property Description
The predefined template is a widely adopted approach(Autili et al., 2015; Dou et al., 2014; Lumpe et al., 2011) to support end-users in defining desired properties. The developer creates templates with placeholders, and the user instantiates them by selecting templates and supplying specific values for placeholders. These templates can be translated into formal language such as linear temporal logic(LTL) and metric temporal logic(MTL), which are widely used for property description and system verification (Clarke et al., 1986).
An LTL formula consists of atomic propositions (basic statements that cannot be further decomposed, i.e., the state value of space and device), logical operators (such as and $\land$, or $\lor$, and implies $\rightarrow$), and temporal operators (such as globally G and finally F). MTL extends LTL by adding time-related metrics to specify time bounds for temporal operators.
For the commonly desired property types in WoT environment (Xiao et al., 2020; Ibrhim et al., 2021; Zhang et al., 2019; Sun et al., 2023), EnvGuard provides eight templates to support users in describing violations by assigning states, events, or actions to placeholders. In A.2 we list the templates, where LTL-type templates describe violated spatial state, and MTL-type templates describe temporal state trace violation.
Although the template approach enables users to customize system properties, understanding the interplay between device services and environmental context remains complex, which brings challenges to configuring the template accurately. To alleviate the complexity, environment-centric property description templates are further proposed (template #1, #2, #5, and #6) to enable user to describe device behaviors by specifying the effects of the device being executed in the environment. By querying actions with the specified effect as well as the corresponding preconditions in the environment representation, the effect can be instantiated into device state and environmental context. Figure 4 demonstrates an example of brightness management where the abstract effect ”Lab.Brightness_Up” is instantiated to the corresponding states of the laboratory light and curtain as well as the environmental context. The detailed algorithm for instantiating the abstract template is shown in the A.2.
@@
the abstract effect ”Lab.Brightness_Up” is instantiated to the corresponding states of the laboratory light and curtain as well as the environmental context

### 16 Dec 2023 v2

# EnvGuard: Guaranteeing Environment-Centric Safety and Security Properties in Web of Things

@@ Abstract.
Web of Things (WoT) technology facilitates the standardized integration of IoT devices ubiquitously deployed in daily environments, promoting diverse WoT applications to automatically sense and regulate the environment. In WoT environment, heterogeneous applications, user activities, and environment changes collectively influence device behaviors, posing risks of unexpected violations of safety and security properties. Existing work on violation identification primarily focuses on the analysis of automated applications, lacking consideration of the intricate interactions in the environment. Moreover, users’ intention for violation resolving strategy is much less investigated. To address these limitations, we introduce EnvGuard, an environment-centric approach for property customizing, violation identification and resolution execution in WoT environment. We evaluated EnvGuard in two typical WoT environments. By conducting user studies and analyzing collected real-world environment data, we assess the performance of EnvGuard, and construct a dataset from the collected data to support environment-level violation identification. The results demonstrate the superiority of EnvGuard compared to previous state-of-the-art work, and confirm its usability, feasibility and runtime efficiency.
Web of Things, violation identification and resolving, dataset
@@ 1. Introduction
The rapid development of Internet of Things (IoT) has accelerated the deployment of smart devices in daily environments. Building on this, Web of Things (WoT) standardizes the integration of fragmented IoT device as perceptual event and executable action service using web technologies (Sun et al., 2023; Zyrianoff et al., 2022), enhancing device interoperability (Lu et al., 2023; Sharma et al., 2017) and facilitating the development of various WoT applications for automated sensing and regulation of environments (Guinard et al., 2011).
Although the WoT paradigm brings great convenience for daily life, safety and security concerns attract great attention (Pradeep et al., 2021; Huang et al., 2021; Chaki et al., 2020; Chaki and Bouguettaya, 2020; Wang et al., 2019; Shah et al., 2019). It is crucial to guarantee the device behavior complies with the desired environment properties, i.e., the safety and security constraints that should always be satisfied by the WoT environment. For example, a common safety concern is the improper use of electrical devices like air conditioners or lights, posing threats to people’s health and the electrical safety standards. Another security concern is the risk of property loss and privacy breach due to windows and curtains remaining open when the house is empty.
To this end, existing research contributes to identifying property violations by analyzing the impact of WoT applications on the state of devices, and resolves them by intercepting or disabling applications (Çelik et al., 2018; Nguyen et al., 2018; Hsu et al., 2019; Chen et al., 2019; Chi et al., 2018; Alhanahnah et al., 2023; Ding et al., 2021; Zhang et al., 2019). However, limitations still remain in many aspects.
Property violations are not only caused by applications, but also by user activities and environment changes which current application-centric methods fail to identify. For example, a user manually opens the meeting room window on a sunny day and leaves, and subsequently the weather changes to rain, posing a risk of rainwater damage to indoor items.
@@
The interplay between environment context and the context-sensitive device services implies that current methods lack analysis of the intricate interrelation. For example, opening and closing a window can have varying effects on indoor temperature, humidity, and air quality under different outdoor conditions. Identifying violations only based on the application-related device state can lead to omissions or misjudgments.
@@
The definition of property fails to meet the customization requirements when it is predefined and built-in, while accurately specifying properties is challenging for non-technical users. For example, to specify properties related to the environment temperature, users must consider all devices potentially affecting temperature (e.g., air conditioner, window, and heater) and the related environment context (e.g., weather, outdoor temperature).
@@
Environment properties span both spatial and temporal aspects, while existing methods primarily address spatial environment context, neglecting the temporal trace of device services. For instance, ensuring electrical safety by detecting prolonged lighting in empty rooms requires a temporal analysis of human detection events and light-switch actions.
@@
The intercepting or revoking resolving strategy may be inconsistent with user intention, as the violation actions possibly contain user requirements. For example, if a user opens a window while raining, closing it as a resolution might be contrary to the ventilation intent.
To address these limitations, we present EnvGuard, an environment-centric approach that offers expressive descriptions and comprehensive guarantees for properties, as shown in Figure 1
@@
EnvGuard first defines a conceptual schema to model the interplay between environment context and context-sensitive device services, and automatically generates a representation of a WoT environment based on device and space information. Then, a GUI tool is provided to assist users in customizing spatial and temporal environment properties by instantiating property description templates, which can be translated into linear temporal logic (LTL) and metric temporal logic (MTL) formulas. During offline phase, EnvGuard separately analyzes the LTL describing the spatial state properties and the MTL describing the temporal trajectory properties by adopting a hybrid model-checking technique, identifying potential violations and their possible resolution actions. According to the result of the user study on user preference for resolving different types of violations, three optional resolving strategies are proposed to ensure the consistency of resolution action and user’s intention. At runtime, EnvGuard checks spatial state and temporal trace property in an event-driven manner and, based on the resolving strategy, selects non-conflicting actions to resolve identified violations without introducing new ones.
To evaluate EnvGuard, we implement it as a prototype and conduct two representative real-world case studies including a smart office and a smart home. The evaluation demonstrates that EnvGuard can facilitate users in constructing spatio-temporal environment properties, accurately identifies all property violations, and generates resolution actions aligned with user intentions while ensuring real-time performance.
We summarize our main contributions as follows:
Violation Identification. We introduce a schema to model the interplay between environment context and devices, enabling comprehensive identification of spatio-temporal property violations caused by various entities (including application, user activities, and environment changes) through a hybrid model-checking technique.
Resolving Strategy. We investigate user preferences for resolution actions, offering selectable resolving strategies for different types of properties to ensure alignment with user intentions.
Property Customizing. We provide multiple property description templates and corresponding GUI tool to assist users in customizing environment properties.
Practical Results. We conduct two real-world case studies to demonstrate EnvGuard’s performance, and construct the collected environment data as the first dataset(EnvGuard, 2023) which records events and actions in the environment and manually labels 20 violation types.
@@ 2. Environment Modeling
In this section, we first present a schema to model the interplay between environment context and device service, and introduce its automatic construction based on device and space information.
@@ 2.1. Conceptual Schema of Environment
The conceptual schema is illustrated in Figure 2, describing the concepts and the relationship of the spaces, the events and actions provided by devices, and the effects that actions can produce on the environment under the specific preconditions. We further formalize each concept in the conceptual schema.
@@
Space describes the space within the environment, which is formally defined as a tuple of ¡id, type, Attributes, States¿. id is a unique identifier of the space and type is the category of the space (e.g., meeting room, corridor, etc.). Attributes are a set of space attributes defined as a map $\{(a_{1}:v_{1}),...,(a_{n}:v_{n})\}$ where each $a_{i}$ is the attribute name (e.g., area) and $v_{i}$ is the corresponding value (e.g., 50 $m^{2}$). And States is a set of variables of the space defined as a map $\{(s_{1}:v_{1}),...,(s_{n}:v_{n})\}$ where each $s_{i}$ is the state name (e.g., temperature) and $v_{i}$ is the corresponding value. The continuous state values are discretized (e.g., discretize the temperature into high, medium, and low). The topological structure of different spaces is described by the ”adjacentTo” and ”reachableTo” relationship.
@@
Device describes the deployed device, and is represented as a tuple of ¡id, type, state¿, indicating the unique identifier, category (e.g., window, light), and current state (e.g., on, off). The ”locatedIn” relationship specifies the space where a device is deployed.
@@
Event is a device service that describes changes in the environment state and is provided by sensors. Event is defined as a tuple of ¡event_type, location, target_state, payload, Topic¿. event_type indicates the category of the event (e.g., Temperature_Change), location indicates the space where the event occurs, target_state specifies the space state impacted by the event (e.g., temperature), and payload is the current value of the changed state (e.g., high). Topic is defined as a map $\{(id_{1}:t_{1}),...,(id_{n}:t_{n})\}$ where each $id_{i}$ indicates the id of device that can provide the event, and each $t_{i}$ indicates the message queue topic for subscribing to the event provided by the device. The ”providedBy” relationship specifies the device providing the event, and the ”update” relationship specifies the target space where the state change has occurred.
@@
Action is the other device service that describes the executable function of actuators, which is defined as a tuple of ¡action_type, location, target_state, payload, URL¿. action_type categorizes the action (e.g., AC_TurnOn), location specifies the execution site, target_state indicates the device whose state is impacted by the action, and the payload is the value of the current device state. URL is a map $\{(id_{1}:url_{1}),...,(id_{n}:url_{n})\}$ where each $id_{i}$ is a device ID offering the action, and each $url_{i}$ is the API address for the action service. The ”providedBy” relationship specifies the device performing the action, and the ”has” relationship specifies the action’s impact on environment contexts.
@@
Effect describes the impact of an action on the environment state, which is defined as a pair of ¡state_name, effect_type¿. state_name specifies the space state impacted by the action (e.g., temperature), and effect_type specifies the effect of the action on the target_state (e.g., increase, decrease, etc.). The ”affect” relationship indicates the space impacted by this effect, and the ”constrainedBy” relationship specifies the constraint condition.
PreCondition indicates the prerequisite environment conditions that need to be satisfied for the effect to take place. PreCondition is defined as a tuple of ¡state_name, operator, target_value¿, where state_name specifies the space state that needs to be satisfied, operator specifies the comparison operator (e.g., ¿, ¡, etc.), and target_value specifies the value that needs to be satisfied. Specifically, the target_value can be set either by a specific value or referring to a space state. For example, the target_value can be set to high or refer to the temperature of the adjacent space. The ”referTo” relationship specifies the space that the state_name and target_value refer to.
@@ 2.2. Environment Representation Construction
All instances generated by the device description format can be automatically associated with space instances based on the device’s location information (e.g., the window instance is connected to the Lab01 instance through the ”locatedIn” relationship), completing the construction of environment representation.
@@ 3. Property Description
EnvGuard employs the widely adopted template approach(Autili et al., 2015; Dou et al., 2014; Lumpe et al., 2011) to support end-users in customizing desired environment properties. The users specify environment properties by selecting templates and instantiating them with specific values for placeholders in the templates. These templates can be translated into formal language such as linear temporal logic(LTL) and metric temporal logic(MTL), which are widely used for property description and system verification (Clarke et al., 1986). An LTL formula consists of atomic propositions (basic statements that cannot be further decomposed, i.e., the state value of space and device), logical operators (such as and $\land$, or $\lor$, and implies $\rightarrow$), and temporal operators (such as globally G and finally F). MTL extends LTL by adding time-related metrics to specify time bounds for temporal operators.
For the commonly desired property types in WoT environment (Xiao et al., 2020; Ibrhim et al., 2021; Zhang et al., 2019; Sun et al., 2023), EnvGuard provides eight templates to support users in describing violations by assigning states, events, or actions to placeholders. Table 1 lists the templates, where LTL-type templates describe violated spatial state, and MTL-type templates describe temporal state trace violation.
Table 1. Property description templates and the corresponding LTL/MTL formulas. The superscript ${}^{+}$” indicates that multiple attributes can be specified.
Type Id Property Description Template LTL/MTL Formula
[…]
Although the template approach enables users to customize system properties, understanding the interplay between device services and environment context remains complex, which brings challenges to configuring the template accurately. To alleviate the complexity, environment-centric property description templates are further proposed (template #1, #2, #5, and #6) to enable user to describe device behaviors by specifying the effects of the device being executed in the environment. By querying actions with the specified effect as well as the corresponding preconditions in the environment representation, the effect can be instantiated into device state and environment context. Figure 4 demonstrates an example of brightness management where the abstract effect ”Lab.Brightness_Up” is instantiated to the corresponding states of the laboratory light and curtain as well as the environment context.
@@
the abstract effect ”Lab.Brightness_Up” is instantiated to the corresponding states of the laboratory light and curtain as well as the environment context